OpenStudy (anonymous):

how do i solve 4e^-2x=17?

OpenStudy (inkyvoyd):

4e^-2x=17 or 4e^(-2x)=17?

OpenStudy (anonymous):

4e^(-2x)=17

OpenStudy (inkyvoyd):

divide both sides by 4 to get \(\Huge e^{-2x}=17/4\) take the natural log of both sides to get -2x=ln(17/4) then solve for x

OpenStudy (inkyvoyd):

I'll leave the last step(s) to you :)

OpenStudy (anonymous):

thank you!